OpenStudy (emilyjones284):

Ill give a medal (: What is the solution of the equation 3^x =12?

OpenStudy (souvik):

take log in both sides

OpenStudy (emilyjones284):

I dontnw how to do thatok

OpenStudy (emilyjones284):

Sorry my keyboard is messing up

OpenStudy (souvik):

do u know logarithms ?

OpenStudy (emilyjones284):

!no

OpenStudy (souvik):

there is property of logarithms that...\[\log_{10}a ^{b}=b \log_{10}a \] if you take log in your equation... \[\log_{10}3^{x}=\log_{10}12 \] now use that property..

OpenStudy (emilyjones284):

Wait i dont understand what i would do after that

OpenStudy (anonymous):

The answer is x=\[\frac{ \log_{12} }{\log_{3} ? }\]

OpenStudy (jdoe0001):

$$\large { 3^x =12\\ log_3(3^x) = log_3(12)\\ \text{using the log cancellation rule of}\\ \color{blue}{log_aa^x = x}\\ x = log_3(12) } $$

OpenStudy (emilyjones284):

Ohh a o(: uy knah tdnatsrednuiykohh

OpenStudy (jdoe0001):

and you can use the "change of base rule" to get the actual value using \(log_{10}\) only $$ large { log_3(12)\\ \text{using log change of base rule}\\ log_3(12) \implies \cfrac{log_{10}12}{log_{10}3} } $$

OpenStudy (jdoe0001):

woops, darnheh

OpenStudy (jdoe0001):

$$ \large { log_3(12)\\ \text{using log change of base rule}\\ log_3(12) \implies \cfrac{log_{10}12}{log_{10}3} } $$

OpenStudy (emilyjones284):

Oops im sorrym oard is messing up on my ipad